If the sides of a square are decreased by 3 cm, the area is decreased by 81 cm2. What were the dimensions of the original square?
Answer by checkley75(3666)

x^2=(x-3)^2+81
x^2=x^2-6x+9+81
6x=90
x=90/6
x=15 cm. for the original length of the sides of the square.
proof
15^2=(15-3)^2+81
225=12^2+81
225=144+81
225=225